OpenStudy (anonymous):

Will give MEdal and FAn!(:

OpenStudy (anonymous):

@jim_thompson5910 @ryan123345

OpenStudy (anonymous):

\[-10=4(-2)+2\]\[-10=-8+2\]\[-10=-6\] no not the first one

OpenStudy (anonymous):

\[6=4\times 1+2\]\[6=4+2\]\[6=6\] yes for the second one

OpenStudy (ryan123345):

could u pick more than 1 answer sparkles?

OpenStudy (anonymous):

yup! (: @ryan123345

OpenStudy (anonymous):

you have 2 lolz

OpenStudy (ryan123345):